OpenStudy (anonymous):

How do you graph y=(2|x-1|)/x-1?

OpenStudy (korosh23):

Use graphing calculator.

OpenStudy (anonymous):

I know the answer, but I need to know how to graph it myself.

OpenStudy (alekos):

just do it the old fashioned way plug in some values for x and work out the y value x = -5,-4,-3,-2,-1,0,1,2,3,4,5

OpenStudy (anonymous):

Is there a reason why the graph doesn't look like a v, even with absolute value in the equation?

OpenStudy (alekos):

because we are dividing by x-1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

don't be fooled by this, it is just two horizontal lines

OpenStudy (misty1212):

\(|x-1|=1-x\) if \(x<1\) or \(x-1\) if \(x>1\)

OpenStudy (alekos):

\[y = 2\left| x-1 \right|\] this would give us a V shaped graph

OpenStudy (misty1212):

so all you really have is \(-2\) if \(x<1\) or \(2\) if \(x>1\)

OpenStudy (anonymous):

So the y is constant, and it's not in the shape of a v because we are dividing by x-1?

OpenStudy (misty1212):

\[\frac{2(x-1)}{x-1}=2\\ \frac{2(1-x)}{x-1}=-2\]

OpenStudy (misty1212):

in other words, this is a rather silly way of just saying \[f(x) = \left\{\begin{array}{rcc} -2& \text{if} & x <1 \\ 2& \text{if} & x >1 \end{array} \right. \]